# RoboTwin 2.0: A Scalable Data Generator and Benchmark with Strong Domain Randomization for Robust Bimanual Robotic Manipulation

### v1

@@ 6 Conclusion
This paper presented RoboTwin 2.0, a scalable simulation framework for generating diverse, high-fidelity expert data to support robust bimanual manipulation. Our system integrates MLLM-based task generation, embodiment-adaptive behavior synthesis, and comprehensive domain randomization to address key limitations in prior synthetic datasets.
@@ Appendix C Domain Randomization Setting
Domain randomization in all experiments includes cluttered scenes, random lighting, table height variation (up to 3 cm), and randomized background textures.
@@ Appendix D Policies Training Details
RDT was pretrained for 100,000 steps with a batch size of 16 per GPU on 8 GPUs, and subsequently fine-tuned for 10,000 steps with a batch size of 16 per GPU on 4 GPUs.
Pi0 was pretrained for 100,000 steps with a batch size of 32, and fine-tuned for 30,000 steps using the same batch size.
For all ACT-based policies, we use a unified training setup with a chunk size of 50, batch size of 8, and single-GPU training for 6000 steps. During deployment, we apply temporal_agg for temporal aggregation to enhance execution stability.
DP was trained for 600 steps with a batch size of 128 and a planning horizon of 8.
DP3 was trained for 3,000 steps with a batch size of 256, using a planning horizon of 8 and a point cloud resolution of 1,024.
D.1 Task-Specific Performance Comparison on Code Generation
@@
pick_dual_bottles_easy 99% 99% 100% 100% 100% 100%
@@
Appendix E Prompts for Generating Task Instructions and Object Descriptions
@@
Listing 1: Prompts for Generating Task Instructions and Object Descriptions
Appendix F Task Instruction and Object Description Example
Object Description # object id - ‘001_bottle/0’:
# object id - ‘039_mug/0’:
We use the DeepSeek-V3 model for program synthesis and the moonshot-v1-32k-vision-preview model for multimodal error localization and verification. These models were selected for their strong performance in language reasoning and visual understanding while maintaining efficiency suitable for large-scale iterative refinement. The success rate of the $i$-th program is computed as $R_{i}=\frac{1}{M}\sum_{j=1}^{M}s_{i,j}$, and the final success rate for a given task under a specific system variant is then defined as $R_{\text{task}}=\frac{1}{N}\sum_{i=1}^{N}R_{i}$.
G.1 Metric Definitions
We report the following metrics across all tasks:
CRSR (Code Revision Success Rate) measures the final success rate of each task after up to five rounds of code repair, incorporating either execution feedback (FB) or multimodal feedback (MM FB).
Token denotes the average number of tokens generated by the language model per task, including both initial synthesis and all revision steps. It serves as a proxy for computational cost and LLM inference budget.
Appendix H Success Rates of Different Embodiments on RoboTwin 2.0 Tasks
We report the success rates of five robot embodiments across the 50 RoboTwin 2.0 tasks, using the same set of expert code for data generation, as shown in Table 9.
@@ G.1 Metric Definitions
Pick Dual Bottles Easy 8% 3% 0% 0% 0% 92% 6% 0% 81% 7%

### v2

@@ 6 Conclusion
This paper presented RoboTwin 2.0, a scalable simulation framework for generating diverse, high-fidelity expert data to support robust bimanual manipulation. Our system integrates MLLM-based task generation, embodiment-adaptive behavior synthesis, and comprehensive domain randomization to address key limitations in prior synthetic data generator.
@@ Appendix C Domain Randomization Setting
Domain randomization in all experiments includes cluttered scenes, random lighting, table height variation (up to 3 cm), unseen language instructions and randomized background textures.
@@ Appendix D Policies Training Details
RDT in experiment 4.3 was pretrained for 100,000 steps with a batch size of 16 per GPU on 8 GPUs, and all single-task fine-tuning was conducted for 10,000 steps with a batch size of 16 per GPU on 4 GPUs.
Pi0 in experiment 4.3 was pretrained for 100,000 steps with a batch size of 32, and all fine-tuning was performed for 30,000 steps using the same batch size.
ACT was trained under a unified setup with a chunk size of 50, batch size of 8, and single-GPU training for 6,000 epochs. During deployment, we applied temporal_agg for temporal aggregation to improve execution stability.
DP was trained for 600 epochs with a batch size of 128 and a planning horizon of 8.
DP3 was trained for 3,000 epochs with a batch size of 256, using a planning horizon of 8 and a point cloud resolution of 1,024, with precise segmentation of the background and tabletop.
Appendix E Support for Flexible Embodiment Combinations
Appendix F Improvements of RoboTwin 2.0 over RoboTwin 1.0 Policy Codebase
Metric RoboTwin 1.0 RoboTwin 2.0
Prompt Token Length ↓ 5901.0 4719.1
Code Token Length ↓ 1236.6 569.4
Parallelism Control ↑
AST Similarity wen2019code 23.72% 44.78%
CodeBLEU Similarity ren2020codebleu 17.18% 18.53%
CodeBERT Similarity feng2020codebert 97.72% 98.80%
Unixcoder Similarity guo2022unixcoder 76.24% 82.21%
Avg. VLM Token Cost (per observation) 6894
Table 7: Code Generation Efficiency and Quality Comparison. Evaluation of prompt and generated code characteristics, along with code similarity metrics (AST Structural Similarity, CodeBERT, Unixcoder cosine similarity) against expert-written code, for RoboTwin 1.0 and RoboTwin 2.0 in zero-shot generation. The VLM observer cost is also reported for RoboTwin 2.0.
We first quantify the architectural impact of RoboTwin 2.0 in a one-shot generation without code repair and iterative refinement. Table 7 shows that RoboTwin 2.0 yields significantly shorter programs (569.4 vs. 1236.6 tokens), with reduced prompt length and higher structural similarity to human-written code. Crucially, it enables dual-arm parallelism via a unified API abstraction, which is absent in RoboTwin 1.0.
These improvements stem from the structured prompting and geometric API modularization designed into RoboTwin 2.0. Higher AST similarity (+21.06%), CodeBERT similarity (+1.08%), and Unixcoder alignment (+5.97%) indicate that RoboTwin 2.0 not only reduces code size but also improves semantic clarity and functional alignment.
In addition, RoboTwin 2.0 integrates a VLM observer, a plug-and-play module triggered only when execution fails. To quantify its overhead, we estimated VLM usage via the Kimi API (assuming each image = 1,024 tokens) over three representative tasks: the average cost was 6,295 input tokens and 599 output tokens, totaling 6,894 tokens. While this introduces moderate overhead, the VLM enables RoboTwin 2.0 to catch and correct errors invisible to execution logging, significantly enhancing robustness and overall task success. Importantly, the observer remains optional and can be disabled when prioritizing token efficiency.
We use the DeepSeek-V3 model for program synthesis and the moonshot-v1-32k-vision-preview model for multimodal error localization and verification. These models were selected for their strong performance in language reasoning and visual understanding while maintaining efficiency suitable for large-scale iterative refinement. The success rate of the $i$-th program is computed as $R_{i}=\frac{1}{M}\sum_{j=1}^{M}s_{i,j}$, and the final success rate for a given task under a specific system variant is then defined as $R_{\text{task}}=\frac{1}{N}\sum_{i=1}^{N}R_{i}$. For detailed usage instructions, refer to https://robotwin-platform.github.io/doc/usage/expert-code-gen.html.
G.1 Metric Definitions
We report the following metrics across all tasks:
Token denotes the average number of tokens of policy code generated by the language model per task. It serves as a proxy for computational cost and LLM inference budget.
G.2 Task-Specific Performance Comparison on Code Generation
@@
pick_dual_bottles 99% 99% 100% 100% 100% 100%
@@
G.3 Per-task Success Rates of Code Generation
We report the success rates of all tasks in Tab. 9.
Task Rate Task Rate Task Rate Task Rate
Click Alarmclock 0% Click Bell 10% Dump Bin Bigbin 0% Grab Roller 74%
Handover Block 27% Handover Mic 0% Hanging Mug 0% Lift Pot 40%
Stamp Seal 20% Turn Switch 0% Avg Success Rate 43.34%
G.4 Multimodal Observation and Error Localization
To further investigate the capability of the VLM observer, we manually curated a dataset of 130 execution sequences, including 101 failed trials and 29 successful trials. Each sequence consists of the natural language task instruction, a series of visual observations, and policy code. This dataset enables us to evaluate both binary error detection and fine-grained error localization.
Error Detection.
The VLM observer was first tasked with evaluating whether a robotic execution successfully completed the instructed task. The confusion matrix is as follows: TP = 16, FP = 61, TN = 40, FN = 13. The derived performance metrics are: Accuracy = 0.431, Precision = 0.208, Recall = 0.552, and F1-score = 0.302. These results suggest that while the VLM observer is reasonably sensitive in identifying execution failures (recall), its tendency to over-predict errors leads to a high false positive rate.
Error Localization.
For failed executions, we further tested whether the VLM observer could correctly localize the program step or code snippet that led to failure. Among the 40 true negative samples (correctly identified failed trials), the observer successfully pinpointed the failure step in 12 cases, yielding a localization accuracy of 30%. This indicates that, while the observer shows promising diagnostic ability, there remains significant room for improvement in reliable error attribution.
Case Study.
We illustrate one successful case of VLM observer reasoning. Given a task instruction and five visual observations, the observer produced a step-by-step analysis:
Step 1: The left arm successfully picks up the mug from the left side.
Step 2: The left arm successfully places the mug at the middle position.
Step 3: The right arm successfully picks up the mug from the middle position.
Step 4: The right arm attempts to hang the mug on the rack but fails.
Step 5: The right arm is in the process of moving away after attempting to hang the mug.
The overall task was not successfully completed. The failure occurred in Step 4, where the right arm attempted to hang the mug on the rack but did not succeed. The mug is not properly positioned on the rack. The error message in the code indicates an AttributeError: ’list’ object has no attribute ’p’. This error occurs in the line where the target_point is being set to middle_target_pose.p. The issue is likely due to the fact that middle_target_pose is a list, and lists do not have an attribute ’p’. The code should be modified to correctly access the position data from the middle_target_pose list.
To fix the error, the code should be updated to access the position data correctly. For example, if middle_target_pose is a list containing the position and orientation, the position data should be accessed using middle_target_pose[0] instead of middle_target_pose.p. Additionally, the failure in Step 4 could be due to the incorrect position or orientation of the mug when attempting to hang it on the rack. The code should be reviewed to ensure that the mug is positioned and oriented correctly before executing the hanging motion.
Failure Modes.
Despite these successes, we also observed cases where the VLM observer fails. Misclassifications primarily occur in failed executions: the observer may hallucinate correct outcomes or overlook subtle orientation errors. For instance, in a stacking task where boxes are misaligned, the observer sometimes marks the execution as successful if all boxes are placed on top of each other. Similarly, in tasks requiring fine-grained pose accuracy (e.g., placing shoes in the correct orientation), small but critical errors may be ignored. Furthermore, failures stemming from invisible factors—such as incorrect grasp axis parameters—remain challenging for purely vision-based observers to diagnose.
G.5 LLM-Generated Code and Human-Written Code Case Study
LLM-Generated Code (gpt_place_shoe)
class gpt_place_shoe(place_shoe):
def play_once(self):
# Initial observation
self.save_camera_images(task_name="place_shoe", step_name="step1_initial_scene_state", generate_num_id="generate_num_0")
# Get the shoes position to determine which arm to use
shoe_pose = self.shoe.get_pose()
shoe_position = shoe_pose.p
arm_tag = ArmTag("left" if shoe_position[0] < 0 else "right")
# Grasp the shoe
self.move(self.grasp_actor(actor=self.shoe, arm_tag=arm_tag, pre_grasp_dis=0.1, grasp_dis=0))
self.save_camera_images(task_name="place_shoe", step_name="step2_shoe_grasped", generate_num_id="generate_num_0")
# Lift the shoe up to avoid collision
self.move(self.move_by_displacement(arm_tag=arm_tag, z=0.07, move_axis=’world’))
self.save_camera_images(task_name="place_shoe", step_name="step3_shoe_lifted", generate_num_id="generate_num_0")
# Get target pose from the top functional point of the target block
target_pose = self.target_block.get_functional_point(1, "pose")
# Place the shoe on the target block with head towards left side
self.move(self.place_actor(actor=self.shoe, arm_tag=arm_tag, target_pose=target_pose, functional_point_id=0, pre_dis=0.1, dis=0.02, is_open=True, constrain="align", pre_dis_axis=’fp’))
self.save_camera_images(task_name="place_shoe", step_name="step4_shoe_placed", generate_num_id="generate_num_0")
# Lift the gripper up after placing
self.move(self.move_by_displacement(arm_tag=arm_tag, z=0.07, move_axis=’world’))
# Return arm to origin
self.move(self.back_to_origin(arm_tag=arm_tag))
# Final observation
self.move(self.save_camera_images(task_name="place_shoe", step_name="step5_final_scene_state", generate_num_id="generate_num_0"))
Listing 1: LLM-Generated Code for gpt_place_shoe Task
Human-Written Code (place_shoe)
class place_shoe(base_task):
def play_once(self):
# Get the shoes position to determine which arm to use
shoe_pose = self.shoe.get_pose().p
arm_tag = ArmTag("left" if shoe_pose[0] < 0 else "right")
# Grasp the shoe with specified pre-grasp distance and gripper position
self.move(self.grasp_actor(self.shoe, arm_tag=arm_tag, pre_grasp_dis=0.1, gripper_pos=0))
# Lift the shoe up by 0.07 meters in z-direction
self.move(self.move_by_displacement(arm_tag=arm_tag, z=0.07))
# Get target_blocks functional point as target pose
target_pose = self.target_block.get_functional_point(0)
# Place the shoe on the target_block with alignment constraint and specified pre-placement distance
self.move(self.place_actor(self.shoe, arm_tag=arm_tag, target_pose=target_pose, functional_point_id=0, pre_dis=0.12, constrain="align"))
# Open the gripper to release the shoe
self.move(self.open_gripper(arm_tag=arm_tag))
Listing 2: Human-Written Code for place_shoe Task
The LLM generated code tends to be more verbose, explicitly logging intermediate visual states and detailing parameters (e.g., pre_dis_axis=’fp’, is_open=True), while human-written scripts are more minimal, omitting intermediate steps and favoring compact execution. Despite functional similarity, the structural differences illustrate that MLLM-generated programs are not only executable but emphasize step-by-step clarity, contributing to more robust feedback and repair.
Appendix H Task Instruction and Object Description Example
Object Description # object id - ‘001_bottle/0’:
# object id - ‘039_mug/0’:
Appendix I Prompts for Generating Task Instructions and Object Descriptions
@@
Listing 3: Prompts for Generating Task Instructions and Object Descriptions
Appendix J RoboTwin 2.0 Benchmark Setting Visualization
We visualize the simulation settings of the RoboTwin 2.0 benchmark in Fig. 12. All models are trained on 50 clean (non-randomized) demonstrations per task (blue). For evaluation, the Easy setting also uses clean environments, while the Hard setting employs domain-randomized environments (green).
Appendix K Full RoboTwin 2.0 Benchmark
We report the evaluation results of five policies on the RoboTwin 2.0 benchmark under the Easy and Hard settings. Note that these two settings differ only in evaluation conditions, while the training setup remains identical. A continuously maintained online leaderboard is available at https://robotwin-platform.github.io/leaderboard.
Table 10: RoboTwin 2.0 Simulation Benchmark (clean vs randomized, 50+ tasks).
Simulation Task RDT Pi0 ACT DP DP3
Easy Hard Easy Hard Easy Hard Easy Hard Easy Hard
Adjust Bottle 81% 75% 90% 56% 97% 23% 97% 0% 99% 3%
Beat Block Hammer 77% 37% 43% 21% 56% 3% 42% 0% 72% 8%
Blocks Ranking RGB 3% 0% 19% 5% 1% 0% 0% 0% 3% 0%
Blocks Ranking Size 0% 0% 7% 1% 0% 0% 1% 0% 2% 0%
Click Alarmclock 61% 12% 63% 11% 32% 4% 61% 5% 77% 14%
Click Bell 80% 9% 44% 3% 58% 3% 54% 0% 90% 0%
Dump Bin Bigbin 64% 32% 83% 24% 68% 1% 49% 0% 85% 53%
Grab Roller 74% 43% 96% 80% 94% 25% 98% 0% 98% 2%
Handover Block 45% 14% 45% 8% 42% 0% 10% 0% 70% 0%
Handover Mic 90% 31% 98% 13% 85% 0% 53% 0% 100% 3%
Hanging Mug 23% 16% 11% 3% 7% 0% 8% 0% 17% 1%
Lift Pot 72% 9% 84% 36% 88% 0% 39% 0% 97% 0%
Move Can Pot 25% 12% 58% 21% 22% 4% 39% 0% 70% 6%
Move Pillbottle Pad 8% 0% 21% 1% 0% 0% 1% 0% 41% 0%
Move Playingcard Away 43% 11% 53% 22% 36% 0% 47% 0% 68% 3%
Move Stapler Pad 2% 0% 0% 2% 0% 0% 1% 0% 12% 0%
Open Laptop 59% 32% 85% 46% 56% 0% 49% 0% 82% 7%
Open Microwave 37% 20% 80% 50% 86% 0% 5% 0% 61% 22%
Pick Diverse Bottles 2% 0% 27% 6% 7% 0% 6% 0% 52% 1%
Pick Dual Bottles 42% 13% 57% 12% 31% 0% 24% 0% 60% 1%
Place A2B Left 3% 1% 31% 1% 1% 0% 2% 0% 46% 2%
Place A2B Right 1% 1% 27% 6% 0% 0% 13% 0% 49% 0%
Place Bread Basket 10% 2% 17% 4% 6% 0% 14% 0% 26% 1%
Place Bread Skillet 5% 1% 23% 1% 7% 0% 11% 0% 19% 0%
Place Burger Fries 50% 27% 80% 4% 49% 0% 72% 0% 72% 18%
Place Can Basket 19% 6% 41% 5% 1% 0% 18% 0% 67% 2%
Place Cans Plasticbox 6% 5% 34% 2% 16% 0% 40% 0% 48% 3%
Place Container Plate 78% 17% 88% 45% 72% 1% 41% 0% 86% 1%
Place Dual Shoes 4% 4% 15% 0% 9% 0% 8% 0% 13% 0%
Place Empty Cup 56% 7% 37% 11% 61% 0% 37% 0% 65% 1%
Place Fan 12% 2% 20% 10% 1% 0% 3% 0% 36% 1%
Place Mouse Pad 1% 0% 7% 1% 0% 0% 0% 0% 4% 1%
Place Object Basket 33% 17% 16% 2% 15% 0% 15% 0% 65% 0%
Place Object Scale 1% 0% 10% 0% 0% 0% 1% 0% 15% 0%
Place Object Stand 15% 5% 36% 11% 1% 0% 22% 0% 60% 0%
Place Phone Stand 15% 6% 35% 7% 2% 0% 13% 0% 44% 2%
Place Shoe 35% 7% 28% 6% 5% 0% 23% 0% 58% 2%
Press Stapler 41% 24% 62% 29% 31% 6% 6% 0% 69% 3%
Put Bottles Dustbin 21% 4% 54% 13% 27% 1% 22% 0% 60% 21%
Put Object Cabinet 33% 18% 68% 18% 15% 0% 42% 0% 72% 1%
Rotate QRcode 50% 5% 68% 15% 1% 0% 13% 0% 74% 1%
Scan Object 4% 1% 18% 1% 2% 0% 9% 0% 31% 1%
Shake Bottle Horizontally 84% 51% 99% 51% 63% 4% 59% 18% 100% 25%
Shake Bottle 74% 45% 97% 60% 74% 10% 65% 8% 98% 19%
Stack Blocks Three 2% 0% 17% 0% 0% 0% 0% 0% 1% 0%
Stack Blocks Two 21% 2% 42% 1% 25% 0% 7% 0% 24% 0%
Stack Bowls Three 51% 17% 66% 24% 48% 0% 63% 0% 57% 5%
Stack Bowls Two 76% 30% 91% 41% 82% 0% 61% 0% 83% 6%
Stamp Seal 1% 0% 3% 4% 2% 0% 2% 0% 18% 0%
Turn Switch 35% 15% 27% 23% 5% 2% 36% 1% 46% 8%
Average (%) 34.5 13.7 46.4 16.3 29.7 1.7 28.0 0.6 55.2 5.0
Appendix L Success Rates of Different Embodiments on RoboTwin 2.0 Tasks
Table 11 reports the success rates of five robot embodiments across the 50 RoboTwin 2.0 tasks, using the same set of expert programs for data generation.
@@
Pick Dual Bottles 8% 3% 0% 0% 0% 92% 6% 0% 81% 7%